Determine whether the given ordered pair is a solution of the system of equation. Remember to use alphabetical ordered varibles. (9,1);
8x-y=71
3x-4y=23
Is the ordered pair a solution? No or yes

Answer by checkley77(12844)

(9,1)
8x-y=71
8*9-1=71
72-1=71
71=71
3x-4y=23
3*9-4*1=23
27-4=23
23=23
Answer is YES.